Two organ pipes, closed at one end, when sounded together produce 3 beats/second. If their lengths are in the ratio of 101:100, then the fundamental notes produces by them have the frequencies (in Hz)
(a) 100 and 103
(b) 206 and 203
(c) 300 and 303
(d) 400 and 403

Grade:Upto college level

2 Answers

Arun
3 years ago
The ratio of the lengths of the closed organ pipes is equal to the fundamental  notes produced by them.
That is, L1/L2=f1/f2, since L1/L2=100/101
So in (option D) we have f1/f2=500/505=100/101.
Hence, the fundamental notes (in Hz.) produced by them are 500, 501.
Vikas TU
3 years ago
The ratio of the lengths of the closed organ pipes is equal to the fundamental  notes produced by them.
L1/L1 = f1/f2 
Since L1/L2 = 100/101
So in (option D) we have f1/f2 = 500/505=100/101.
Hence, the fundamental notes (in Hz.) produced by them are 500, 501.